OpenStudy (anonymous):

The line passing through (2,b) and (4,6) is parallel to the line joining (1,4) and (-1,7). Find b

myininaya (myininaya):

So since the line containing the first pair of points is parallel to the line containing the second pair of points Then the lines have the same slope So just find the slope of both lines and set them equal Then solve for b

myininaya (myininaya):

\[\frac{y_1-y_2}{x_1-x_2}=\frac{y'_1-y'_2}{x'_1-x'_2}\] Where \[(x_1=2,y_1=b),(x_2=4,y_2=6) \text{ and } (x'_1=1,y'_1=4), (x'_2=-1,y'_2=7)\]

OpenStudy (anonymous):

okay got it thanks

myininaya (myininaya):

:) cool! :)